OpenStudy (anonymous):

a burgular alarm system has six fail-safe components. the probability of each failing is 0.05. Find these probabilities a)Exactly three will fail b) Fewer than two will fail. c) None will fail. d) Compare the answers from a,b,and c and explain why the results are reasonable

OpenStudy (anonymous):

.05%?

OpenStudy (anonymous):

parden?

OpenStudy (anonymous):

i don't know

OpenStudy (anonymous):

A) 1 in 8000. B) 1 in 20. C) probably wrong on that one, no idea.

OpenStudy (anonymous):

a) P(Exactly 3 fail)= 6C3*0.05^3*0.95^3=0.00214... b)P(Fewer than 2 fail)=P(1 fail)+P(0 fail)=6C1*0.05*0.95^5+0.95^6=0.967... c)P(0 fail)=0.95^6=0.735... Where C denotes the choose function. I think this is right.